# Sharpness-Aware Minimization Can Hallucinate Minimizers

### v1

@@ Abstract
Sharpness-Aware Minimization (SAM) is a widely used method that steers training toward flatter minimizers, which typically generalize better. In this work, however, we show that SAM can converge to hallucinated minimizers—points that are not minimizers of the original objective. We theoretically prove the existence of such hallucinated minimizers and establish conditions for local convergence to them. We further provide empirical evidence demonstrating that SAM can indeed converge to these points in practice. Finally, we propose a simple yet effective remedy for avoiding hallucinated minimizers.
@@
Figure 1: Illustrative example of hallucinated minimizers. See Appendix D for details. (a) Smooth function $f$ with a minimizer set and an isolated maximizer. (b) $f^{\text{SAM}}=f(x+\rho\tfrac{\nabla f(x)}{\|\nabla f(x)\|})$; its minimizers do not correspond to minimizers or stationary points of $f$ and are therefore hallucinated. (c) Vector field of $\nabla f(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|})$; the hallucinated minimizers are attractors of the SAM iteration.
@@ 1 Introduction
It has been empirically observed in deep learning that flat minimizers tend to generalize better than sharp ones (Neyshabur et al., 2017; Jiang et al., 2020). Motivated by this observation, Sharpness-Aware Minimization (SAM) was proposed as a training method that fits the data while simultaneously regularizing against the sharpness (Foret et al., 2021; Zheng et al., 2021; Wu et al., 2020). Specifically, for a differentiable function $f\colon\mathbb{R}^{d}\to\mathbb{R}$, SAM minimizes the objective
$f^{\mathrm{SAM}}(x):=f\!\left(x+\rho\,\frac{\nabla f(x)}{\|\nabla f(x)\|}\right),$
where the perturbation radius $\rho>0$ controls the strength of the regularization. By encouraging convergence to flat minimizers, SAM has demonstrated strong empirical performance and has inspired a variety of practical variants (Foret et al., 2021).
However, although prior theoretical studies have analyzed the convergence of SAM to a minimizer under various conditions, most rely on assumptions that rarely hold in deep learning, such as convexity (Si & Yun, 2024) or a decaying perturbation radius (Khanh et al., 2024; Oikonomou & Loizou, 2025). Consequently, the validity of these convergence results is not guaranteed in practical settings.
@@
In this paper, we theoretically and empirically demonstrate that SAM can, in fact, hallucinate minimizers, in the sense that its iterates may converge to points that are not minimizers of the original objective, as illustrated in Figure 1. This finding reveals a previously unrecognized failure mode of SAM in deep learning, one that is fundamentally distinct from issues caused by local minima or saddle points. To address this, we present a simple switching strategy that effectively avoids hallucinated minimizers.
@@ 1.1 Related work
SAM and its variants.
Building on the observation that flat minimizers are stable under small perturbations (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017; Chaudhari et al., 2019), SAM was proposed as a method to seek such minimizers (Foret et al., 2021). SAM has demonstrated remarkable performance across a wide range of deep learning tasks (Foret et al., 2021; Bahri et al., 2022; Zhong et al., 2022; Lee et al., 2023), motivating numerous extensions.
One line of work improves SAM by modifying its perturbation direction. Kwon et al. (2021) adapted the perturbation in a scale-invariant manner, while Kim et al. (2022) redefined it using Fisher information geometry. Li et al. (2024) removed the full-batch gradient from the perturbation direction and leveraged stochastic gradient noise for generalization. Li & Giannakis (2024) incorporated momentum into the perturbation to suppress variance and stabilize the adversary. Instead of altering the perturbation itself, Zhuang et al. (2022) adjusted the gradient update through orthogonal decomposition to reduce the surrogate gap.
Another line of research addresses the computational overhead of SAM, which stems from requiring two gradient computations per step. Some approaches reduce the number of gradient evaluations: Liu et al. (2022) computed the perturbation only periodically, Jiang et al. (2023) activated SAM only when the gradient norm is large, and Du et al. (2022b) reused gradients to avoid the second computation. Others restrict the scope of perturbations: Du et al. (2022a) applied them only to a random subset of parameters and sharpness-sensitive data, while Mueller et al. (2023) showed that limiting them to normalization layers preserves most of the benefits. Beyond reducing computation, Xie et al. (2024) improved training efficiency by parallelizing the two gradient computations.
@@
Alongside its practical success, a growing body of theoretical work has analyzed SAM and examined its generalization properties from multiple perspectives. Wen et al. (2022) formalized the precise notion of sharpness minimized by SAM, clarifying its regularization effect. Möllenhoff & Khan (2023) reinterpreted SAM as a relaxation of Bayesian inference. Chen et al. (2023) showed that SAM mitigates noise fitting and improves generalization over stochastic gradient descent. Wei et al. (2023) further demonstrated that SAM alone can enhance adversarial robustness while maintaining clean accuracy.
Another line of work investigates the training dynamics and stability of SAM. Compagnoni et al. (2023) analyzed SAM through the lens of stochastic differential equations, offering a continuous-time perspective. Bartlett et al. (2023) studied quadratic objectives, showing how SAM oscillates across narrow valleys before drifting toward wider minimizers. Dai et al. (2023) demonstrated that normalization plays a key role in stabilizing SAM and ensuring robustness. Long & Bartlett (2024) extended the edge-of-stability threshold of gradient descent to SAM, showing that it depends on the gradient norm. More recently, Zhou et al. (2025) highlighted a late-phase effect, whereby SAM selects flatter minimizers when applied in the later stages of training.
Finally, several works have investigated the convergence properties of SAM in diverse settings. Andriushchenko & Flammarion (2022) proposed USAM, an unnormalized variant obtained by removing gradient normalization, and analyzed its convergence. Si & Yun (2024) provided a systematic study across convex, strongly convex, and nonconvex regimes. Khanh et al. (2024) developed a convergence analysis of SAM and its variants within the framework of inexact gradient descent. Most recently, Oikonomou & Loizou (2025) analyzed SAM and USAM within a unified framework and proved convergence under the Polyak–Łojasiewicz condition.
Hallucinated minimizers and our contribution.
Several prior studies have reported that SAM may converge to points that are not minimizers of the original loss. Kaddour et al. (2022) empirically observed that SAM can become trapped at saddle points, and Kim et al. (2023); Compagnoni et al. (2023) provided a theoretical explanation of this phenomenon using a continuous-time model of SAM. In another line of work, Bartlett et al. (2023) showed that the SAM update is equivalent to gradient descent on a surrogate function in the quadratic case, and Si & Yun (2024) proposed a virtual loss to extend this idea, although it is rigorously defined only in one dimension and lacks guarantees in higher dimensions.
In contrast, hallucinated minimizers represent a fundamentally different failure mode of SAM. They differ from saddle points in that they are not critical points of the original loss, and from surrogate-based interpretations in that they arise directly from the SAM objective. Our analysis applies to general nonconvex and high-dimensional settings and provides a rigorous characterization of hallucinated minimizers that can emerge in practical deep learning scenarios.
1.2 Preliminaries and notation
Throughout the paper, we denote $u(x):=\nabla f(x)/\|\nabla f(x)\|$ (for $x$ such that $\nabla f(x)\neq 0$). To optimize the SAM objective $f^{\mathrm{SAM}}(x)$, we require its gradient $\nabla f^{\mathrm{SAM}}(x)$. Under the standard smoothness assumption on $f$, this gradient is given by
$\nabla f^{\mathrm{SAM}}(x)=\big(I+\rho\,\nabla u(x)\big)\,\nabla f\big(x+\rho\,u(x)\big),$
where $\nabla u(x)$ denotes the Jacobian of $u(x)$. In practice, however, for computational simplicity, one does not use the exact gradient $\nabla f^{\mathrm{SAM}}$. Instead, SAM employs the “shifted” gradient $\nabla f\big(x+\rho\,u(x)\big)$. This yields the (full-batch) SAM iteration
$x_{k+1}=x_{k}-\eta_{k}\,\nabla f\big(x_{k}+\rho\,u(x_{k})\big),\qquad k=0,1,2,\dots,$
where $x_{0}\in\mathbb{R}^{d}$ is the starting point and $\eta_{0},\eta_{1},\ldots\in\mathbb{R}_{+}$ is the sequence of step sizes. The perturbation radius $\rho>0$ controls the degree of “flatness”: larger values of $\rho$ expand the neighborhood over which the loss is minimized, thereby steering the SAM iteration towards flatter minimizers. In practice, SAM is typically implemented with stochastic gradients. When clarification is needed, we refer to SAM with full-batch gradients versus SAM with stochastic gradients.
@@ 1.2 Preliminaries and notation
We introduce some notation. A function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is called real-analytic if its Taylor series at any point $x_{0}$ converges to $f$ on a neighborhood of $x_{0}$. For $\alpha\in\mathbb{R}$, the $\alpha$-superlevel set of $f$ is defined as $\{x:\,f(x)\geq\alpha\}$. For a set $C\subset\mathbb{R}^{d}$, we write $\partial C$ for its boundary. A set $C$ is connected if it cannot be expressed as the union of two disjoint, nonempty open sets. The distance from a point $x\in\mathbb{R}^{d}$ to a nonempty set $C\subseteq\mathbb{R}^{d}$ is $d(x,C):=\inf_{y\in C}\|x-y\|$; if $C$ is closed, the infimum is attained, and hence $d(x,C)=\min_{y\in C}\|x-y\|$. For $\delta>0$, $B_{\delta}(x)$ is the ball centered at $x$ with radius $\delta$ and the (closed) $\delta$-neighborhood of $C$ is $\mathcal{N}_{\delta}(C)=\{x:\,d(x,C)\leq\delta\}$.
@@ 2 Existence of hallucinated minimizers
In this section, we establish—under very mild assumptions—the existence of hallucinated minimizers: local minimizers of $f^{\mathrm{SAM}}$ that are not even stationary points of the original function $f$. Formally, we define hallucinated minimizers as follows:
Definition.
A point $x\in\mathbb{R}^{d}$ is a hallucinated minimizer of $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ for $\rho>0$ if $x$ is a local minimizer of $f^{\mathrm{SAM}}=f\Big(x+\rho\tfrac{\nabla f(x)}{\|\nabla f(x)\|}\Big)$ while satisfying $\nabla f(x)\neq 0$.
When the loss function $f$ is convex, hallucinated minimizers cannot arise, as shown in Theorem A.1 of Appendix A. In practical deep learning, however, the loss function $f$ is highly nonconvex, and local maximizers of $f$ can give rise to hallucinated minimizers.
2.1 Simplified existence proof with isolated maximizers
We begin with a simplified proof under the more restrictive assumption that $f$ has an isolated local maximizer, defined as a point $x^{\bullet}$ with an open neighborhood $U$ such that
$\nabla f(x)\neq 0\,\,\text{ and }\,\,f(x)<f(x^{\bullet})\quad\text{for all }x\in U\setminus\{x^{\bullet}\}.$
Theorem 2.1.
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be continuously differentiable. Assume $f$ has a global minimizer (not necessarily unique) and an isolated local maximizer. Then, a hallucinated minimizer exists for some $\rho>0$.

### v2

@@ Abstract
Sharpness-Aware Minimization (SAM) is widely used to seek flatter minima—often linked to better generalization. In its standard implementation, SAM updates the current iterate using the loss gradient evaluated at a point perturbed by distance $\rho$ along the normalized gradient direction. We show that, for some choices of $\rho$, SAM can stall at points where this shifted (perturbed-point) gradient vanishes despite a nonzero original gradient, and therefore, they are not stationary points of the original loss. We call these points hallucinated minimizers, prove their existence under simple nonconvex landscape conditions (e.g., the presence of a local minimizer and a local maximizer), and establish sufficient conditions for local convergence of the SAM iterates to them. We corroborate this failure mode in neural network training and observe that it aligns with SAM’s performance degradation often seen at large $\rho$. Finally, as a practical safeguard, we find that a short initial SGD warm-start before enabling SAM mitigates this failure mode and reduces sensitivity to the choice of $\rho$.
@@
Figure 1: Illustrative example of hallucinated minimizers. See Appendix D for details. (a) Smooth function $f$ with a minimizer set and an isolated maximizer. (b) $f^{\mathrm{SAM}}=f(x+\rho\,u(x))$; its minimizers (labeled as “SAM minimizers” in the plot) do not correspond to minimizers or stationary points of $f$ and are therefore hallucinated. (c) Vector field of the SAM gradient $\nabla f(x^{+})$; the hallucinated minimizers are attractors of the SAM iteration.
@@ 1 Introduction
Sharpness-Aware Minimization (SAM) is widely used to bias training toward flatter regions of the loss landscape (Foret et al., 2021), often correlated with improved generalization (Neyshabur et al., 2017; Jiang et al., 2020). For a differentiable objective $f:\mathbb{R}^{d}\to\mathbb{R}$, SAM is commonly motivated by the min–max formulation
$\min_{x}\max_{\|\epsilon\|\leq\rho}f(x+\epsilon),$
and the inner maximization is typically approximated by a first-order perturbation along the normalized gradient direction. This leads to the surrogate (“SAM”) loss
$f^{\mathrm{SAM}}(x):=f\!\left(x+\rho\,u(x)\right),\qquad u(x)=\frac{\nabla f(x)}{\|\nabla f(x)\|},$
where $\rho>0$ is the perturbation radius controlling the strength of the sharpness regularization.
A practical mismatch in SAM updates.
A central subtlety—often glossed over in high-level descriptions of SAM—is that the standard implementation does not follow the surrogate gradient $\nabla f^{\mathrm{SAM}}(x)$. Instead, it updates using the shifted gradient evaluated at the perturbed point:
$x_{k+1}=x_{k}-\eta_{k}\,\nabla f(x_{k}^{+}),\qquad x^{+}:=x+\rho\,u(x).$
Consequently, the practical SAM dynamics can become stationary whenever $\nabla f(x^{+})=0$, even if $\nabla f(x)\neq 0$. This mismatch becomes especially relevant in the large-$\rho$ regime: in practice, $\rho$ is tuned by sweeps or grid search, and prior work reports performance degradation and instability at larger radii on standard benchmarks (Andriushchenko and Flammarion, 2022). On the theory side, most convergence guarantees assume $\rho$ is below a threshold under various regularity conditions, leaving the large-$\rho$ regime comparatively less understood (Si and Yun, 2023; Khanh et al., 2024; Oikonomou and Loizou, 2025).
Hallucinated minimizers.
We show that the shifted-gradient dynamics can admit a qualitatively distinct failure mode: SAM can stall at points where $\nabla f(x^{+})=0$ despite $\nabla f(x)\neq 0$. We call such points hallucinated minimizers if, in addition, they are local minimizers of the surrogate loss $f^{\mathrm{SAM}}$. Since $\nabla f(x)\neq 0$, hallucinated minimizers are not stationary points of the original objective $f$ (Figure 1). This failure mode is inherently nonconvex (it cannot occur for convex $f$), and becomes particularly relevant in the large-$\rho$ regime, where the empirical performance of SAM is often observed to degrade (Andriushchenko and Flammarion, 2022).
Our results.
We provide a geometric and dynamical characterization of hallucinated minimizers. Specifically, we prove that hallucinated minimizers exist for a nontrivial interval of radii, including values in a large-$\rho$ regime (Theorem 2.2) under simple nonconvex landscape conditions (e.g., the coexistence of a local minimizer and a local maximizer). We extend this existence result from isolated maximizers to bounded local maximizer sets under a real-analyticity assumption (Theorem 2.4). We further show that hallucinated minimizers are not merely an algebraic possibility: an isolated connected component can be locally attracting for the discrete-time SAM dynamics (Theorem 3.1). Finally, we show that hallucinated minimizers can be non-isolated and inherit the dimension of minimizer manifolds via local invertibility of the perturbation map (Theorem 3.2).
Empirically, we provide evidence consistent with this failure mode on neural network objectives using diagnostics that compare shifted vs. unshifted gradients and losses: in smooth full-batch settings we directly verify convergence to points with nearly vanishing shifted gradient but nonzero original gradient, and we observe similar mismatch signatures in larger networks beyond the smooth/analytic setting. Finally, as a practical safeguard, we report that a short initial SGD warm-start before enabling SAM mitigates this failure mode and reduces sensitivity to the choice of $\rho$. Our formal results concern deterministic (full-batch) dynamics; stochastic experiments are included only as empirical observations.
@@
  • We identify and formalize hallucinated minimizers: points where the shifted gradient vanishes, $\nabla f(x^{+})=0$, even though the original gradient is nonzero, $\nabla f(x)\neq 0$, and $x$ is a local minimizer of $f^{\mathrm{SAM}}(x)=f(x^{+})$.
  • We prove existence of hallucinated minimizers for a nontrivial interval of radii (including a large-$\rho$ regime) under mild nonconvex landscape conditions, and extend the result to bounded local maximizer sets under real-analyticity.
  • We provide sufficient conditions under which an isolated hallucinated-minimizer component is locally attracting for the discrete-time SAM dynamics.
  • We show that hallucinated minimizers can form smooth manifolds inheriting the dimension of minimizer manifolds via local invertibility of the perturbation map.
  • We provide empirical diagnostics and observations in neural network training that are consistent with the proposed mechanism.
@@ 1.1 Related work
SAM and practical variants.
SAM was introduced as a method to regularize sharpness by adversarially perturbing parameters during training (Foret et al., 2021). A large body of follow-up work proposes variants that modify the perturbation geometry or the update rule, e.g., scale-invariant perturbations (ASAM) (Kwon et al., 2021), Fisher-geometry perturbations (Kim et al., 2022), incorporation of momentum into the perturbation (Li and Giannakis, 2023) or modified updates such as orthogonal decomposition (GSAM) (Zhuang et al., 2022). Other works reduce SAM’s two-backward-pass overhead via approximation, partial perturbations, or scheduling (e.g., (Liu et al., 2022; Jiang et al., 2023; Du et al., 2022a, b; Mueller et al., 2023)). In this paper, we focus on the standard SAM update—the shifted-gradient step used in common implementations—and analyze its behavior in the large-$\rho$ regime.
@@
Existing theory has examined SAM from multiple perspectives, including sharpness notions and generalization interpretations (Wen et al., 2023; Möllenhoff and Khan, 2023; Chen et al., 2023; Wei et al., 2023). Another line studies SAM’s dynamics and stability, e.g., via continuous-time and SDE approximations (Compagnoni et al., 2023), quadratic dynamics and surrogate interpretations (Bartlett et al., 2023; Si and Yun, 2023), and stability phenomena related to normalization and edge-of-stability effects (Dai et al., 2023; Long and Bartlett, 2024). Several works provide convergence guarantees for SAM and related methods under various assumptions, typically requiring a small-radius condition or a regularity condition that effectively bounds the perturbation effect (e.g., convex/strongly convex regimes, inexact gradient descent frameworks, or PL-type conditions) (Andriushchenko and Flammarion, 2022; Si and Yun, 2023; Khanh et al., 2024; Oikonomou and Loizou, 2025). In contrast, our results characterize a distinct phenomenon that can arise at large $\rho$: the standard SAM dynamics can stall at points that are not stationary (and hence not locally optimal) for $f$.
Differences from prior work: hallucinated minimizers vs. saddle trapping.
Prior work has reported that SAM can become trapped near saddle points (Kaddour et al., 2022), and theoretical explanations have been developed via continuous-time SAM models (Kim et al., 2023; Compagnoni et al., 2023). Hallucinated minimizers represent a qualitatively different failure mode: saddle trapping corresponds to $\nabla f(x)=0$ (stationarity of the original loss), whereas hallucinated minimizers satisfy $\nabla f(x)\neq 0$ but $\nabla f(x^{+})=0$, so they are stationary only for the SAM’s shifted-gradient update. Moreover, our analysis provides $(i)$ geometric existence mechanisms tied to nonconvex landscapes and large $\rho$, $(ii)$ sufficient conditions for local attraction under the discrete-time SAM dynamics, and $(iii)$ manifold structure results via local invertibility of the perturbation map.
Switching schedules.
Switching schedules for SAM have been explored in prior work, often motivated by computational overhead or late-phase effects (Andriushchenko and Flammarion, 2022; Zhou et al., 2025). Complementary to these perspectives, we report that a short early SGD warm-start before enabling SAM can mitigate the hallucinated-minimizer failure mode identified in this paper. Our goal is not to propose a new training recipe, but to provide a simple and broadly compatible safeguard aligned with our mechanism.
1.2 Notation and terminology
@@
We use $\|\cdot\|$ for the Euclidean norm (and its induced operator norm for matrices), write $\mathrm{Sym}(A):=(A+A^{\top})/2$, and let $\lambda_{\min}(A)$ denote the smallest eigenvalue of a symmetric matrix $A$. A function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is called real-analytic if its Taylor series at any point $x_{0}$ converges to $f$ on a neighborhood of $x_{0}$. For $\alpha\in\mathbb{R}$, the $\alpha$-superlevel set of $f$ is defined as $\{x:\,f(x)\geq\alpha\}$. For a set $C\subset\mathbb{R}^{d}$, we write $\partial C$ for its boundary, and we say that $C$ is connected if it cannot be expressed as the union of two disjoint, nonempty open sets. The distance from a point $x\in\mathbb{R}^{d}$ to a nonempty set $C\subseteq\mathbb{R}^{d}$ is $d(x,C):=\inf_{y\in C}\|x-y\|$; if $C$ is closed, the infimum is attained, and hence $d(x,C)=\min_{y\in C}\|x-y\|$. For $\delta>0$, $B_{\delta}(x)$ is the open ball centered at $x$ with radius $\delta$ and $\mathcal{N}_{\delta}(C)=\{x:\,d(x,C)\leq\delta\}$ is the (closed) $\delta$-neighborhood of $C$.
@@ 2 Existence of hallucinated minimizers
In this section, under simple nonconvex landscape assumptions, we establish the existence of hallucinated minimizers: points where the (full-batch) standard SAM update can become stationary even though the original gradient is nonzero.
Why a mismatch is possible.
Recall that SAM is commonly motivated by the surrogate objective
$f^{\mathrm{SAM}}(x):=f\!\left(x+\rho\,u(x)\right),\quad u(x):=\frac{\nabla f(x)}{\|\nabla f(x)\|}$
for $\nabla f(x)\neq 0$. However, the standard SAM implementation does not follow the surrogate gradient $\nabla f^{\mathrm{SAM}}(x)$. Instead, it uses the shifted gradient $\nabla f(x+\rho u(x))$, which we refer to as the SAM gradient. For brevity, we write
$x^{+}:=x+\rho\,u(x).$
Then the (full-batch) SAM iteration takes the form
$x_{k+1}=x_{k}-\eta_{k}\,\nabla f(x_{k}^{+}),$ (1)
so the dynamics can stall whenever $\nabla f(x^{+})=0$, regardless of whether $\nabla f(x)=0$.
Definition 2.1 (Hallucinated minimizers).
Fix $\rho>0$. A point $x\in\mathbb{R}^{d}$ is said to be a hallucinated minimizer at radius $\rho$ if
$\nabla f(x^{+})=0\quad\text{and}\quad\nabla f(x)\neq 0,$
and $x$ is a local minimizer of $f^{\mathrm{SAM}}(x)=f(x^{+})$.
A useful relation between $\nabla f^{\mathrm{SAM}}(x)$ and the SAM gradient $\nabla f(x^{+})$.
Assume in addition that $f$ is twice continuously differentiable on a neighborhood of $x$ and that $\nabla f(x)\neq 0$. Then $u$ is differentiable at $x$ and the chain rule gives
$\nabla f^{\mathrm{SAM}}(x)=\bigl(I+\rho\,\nabla u(x)\bigr)^{\top}\nabla f(x^{+}).$
If $I+\rho\nabla u(x)$ is nonsingular (equivalently, the perturbation map $x\mapsto x^{+}$ is locally invertible at $x$), then $\nabla f^{\mathrm{SAM}}(x)=0$ implies $\nabla f(x^{+})=0$. In particular, at such nondegenerate points, any local minimizer of $f^{\mathrm{SAM}}$ with $\nabla f(x)\neq 0$ automatically satisfies $\nabla f(x^{+})=0$. For brevity, when nondegeneracy is understood, we will therefore sometimes describe hallucinated minimizers simply as local minimizers of $f^{\mathrm{SAM}}$ with $\nabla f(x)\neq 0$. This nondegeneracy condition is used only to justify this shorthand; our existence results do not rely on it.
Convexity rules out hallucinated minimizers.
Hallucinated minimizers cannot arise when $f$ is convex. Indeed, Proposition A.1 in Appendix A shows that if $f$ is convex and $\nabla f(x)\neq 0$, then $\nabla f(x^{+})\neq 0$.
@@
2.1 Simplified existence proof with isolated maximizers
We now present a geometric construction that produces a hallucinated minimizer. For clarity, we begin with the case where $f$ admits an isolated local maximizer, namely, a point $x^{\bullet}$ with an open neighborhood $U$ such that for all $x\in U\setminus\{x^{\bullet}\}$,
$f(x)<f(x^{\bullet})\quad\text{and}\quad\nabla f(x)\neq 0.$
Under this assumption, the following theorem guarantees the existence of a hallucinated minimizer in a large-$\rho$ regime.
Theorem 2.2.
Let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be continuously differentiable. Assume $f$ has a global minimizer $x_{\star}$ (not necessarily unique) and an isolated local maximizer $x^{\bullet}$. Then, there exists a nontrivial interval of radii $\rho$ for which hallucinated minimizers exist, and this interval contains values with $\rho\geq\|x_{\star}-x^{\bullet}\|$.